QUESTION 24

  1. Using the Standard Normal Table. What is the probability a z-score is greater than -0.23?

    In other words, what is P(z > -0.23)?

    A.

    0.9893

    B.

    0.0107

    C.

    0.5910

    D.

    0.4090

    QUESTION 25

  2. Using the Standard Normal Table. What is the probability a z-score is greater than 0.44?

    In other words, what is P(z > 0.44)?

    A.

    0.3300

    B.

    0.3446

    C.

    0.6554

    D.

    0.6700

    QUESTION 26

  3. Using the Standard Normal Table. What is the probability a z-score is between -1.82 and -0.68?

    In other words, what is P( -1.82 < z < -0.68)?

    A.

    0.0422

    B.

    0.2827

    C.

    0.1114

    D.

    0.2139
Answer #1

Solution :

Given that,  

Using standard normal table ,

QUESTION 24

P(z > -0.23) = 1 - P(z < -0.23) = 1 - 0.406 = 0.5910

option c. is correct

QUESTION 25

P(z > 0.44) = 1 - P(z < 0.44) = 1 - 0.67 = 0.3300

option A. is correct

QUESTION 26

P(-1.82 < z < -0.68)

= P(z < -0.68) - P(z < -1.82)

= 0.2483 - 0.0344

= 0.2139

P(-1.82 < z < -0.68) = 0.2139

option D. is correct